# How Many Code and Test Cases Are Enough? Evaluating Test Cases Generation from a Binary-Matrix Perspective

### v1

@@ A.5 Common Failure
Failure to Construct Required Complex Data Structures
Some problems in our benchmark admit only test cases with highly constrained structures. For example, in one problem, every valid input must be a specific type of disconnected graph. However, none of the tested methods successfully produced even a single valid input. As a result, these problems ended up with zero usable test cases. This underscores that generating high-difficulty test cases can be as challenging as solving an algorithmic problem, requiring a deep understanding of both data structures and algorithms.
A.6 Results of Supplementary Experiments
@@ B.1.1 Wrong code
After processing the wrong codes in 2.3, for all retained wrong codes, we used public test cases for testing. For all execution results such as CE, TLE, MLE, EXE, as well as codes that resulted in WA but with empty outputs, manual fixes and reviews were performed. As shown in Figure 9, (a) illustrates a piece of unusable file operation code, in which the script does not include the definitions of Fin and Fout. For this type of code, we removed the corresponding file operations. The error in (b) arises because the unistd.h library already defines a function named link_array, which conflicts with the array link_array defined in the code. (c) presents an example of incomplete code that requires manual supplementation. To ensure consistency between the original and the corrected code, after making modifications we tested the code using private test cases, with the requirement that the test results remain consistent with the crawled results.
@@ B.1.2 Problem Description
Regarding the problem statement processing in 2.3, this subsection provides detailed examples and explanations for problems that heavily rely on images and Special Judge problems.
For problem statements that rely on image-based understanding, such as Stars (see image in reference 10), the problem includes an image that is necessary for understanding in order to generate test cases or solve the problem. We manually reviewed this type of problem statement, filtered out the problems where images affected the understanding of the question, and deleted them. In this step, we deleted a total of 71 problems.
@@
As shown in 13,the following is an example of problem statement cleaning. For demonstration purposes, we have created a sample problem to illustrate the main cleaning tasks. In this process, irrelevant background information is removed, image links and other URLs are discarded, and the phrasing is made smoother. The data range is kept to the most general case. These tasks typically do not follow a universal pattern and require manual inspection. After cleaning the problem statement, we used GPT-4o for translation. In this step, we organized each data entry and deleted 15 problems that were difficult to handle. Then each translation was semantically proofread and certain inappropriate expressions were adjusted for accuracy.The final processed problem statement can be found in B.1.2.

### v2

@@ A.6 Results of Supplementary Experiments
A.7 The Use of Large Language Models (LLMs)
This article utilizes large language models (LLM) solely for writing refinement and graphic enhancement, with no other applications or purposes involved.
@@ B.1.1 Wrong code
After processing the wrong codes in Section2.3, for all retained wrong codes, we used public test cases for testing. For all execution results such as CE, TLE, MLE, EXE, as well as codes that resulted in WA but with empty outputs, manual fixes and reviews were performed. As shown in Figure 11, (a) illustrates a piece of unusable file operation code, in which the script does not include the definitions of Fin and Fout. For this type of code, we removed the corresponding file operations. The error in (b) arises because the unistd.h library already defines a function named link_array, which conflicts with the array link_array defined in the code. (c) presents an example of incomplete code that requires manual supplementation. To ensure consistency between the original and the corrected code, after making modifications we tested the code using private test cases, with the requirement that the test results remain consistent with the crawled results.
@@ B.1.2 Problem Description
Regarding the problem statement processing in Section2.3, this subsection provides detailed examples and explanations for problems that heavily rely on images and Special Judge problems.
For problem statements that rely on image-based understanding, such as Stars (see image in reference Figure12), the problem includes an image that is necessary for understanding in order to generate test cases or solve the problem. We manually reviewed this type of problem statement, filtered out the problems where images affected the understanding of the question, and deleted them. In this step, we deleted a total of 71 problems.
@@
As shown in Figure15,the following is an example of problem statement cleaning. For demonstration purposes, we have created a sample problem to illustrate the main cleaning tasks. In this process, irrelevant background information is removed, image links and other URLs are discarded, and the phrasing is made smoother. The data range is kept to the most general case. These tasks typically do not follow a universal pattern and require manual inspection. After cleaning the problem statement, we used GPT-4o for translation. In this step, we organized each data entry and deleted 15 problems that were difficult to handle. Then each translation was semantically proofread and certain inappropriate expressions were adjusted for accuracy.The final processed problem statement can be found in next page
@@
Appendix C Case Study
To demonstrate the practical effectiveness of our method, we conduct a case study on “Sliding Window”, a classic problem requiring the Monotonic Queue algorithm. The problem involves an integer array of length $N(\leq 10^{6})$ and a window of size $K(\leq 10^{6})$. The window slides from the leftmost to the rightmost of the array, moving one position at a time. The goal is to determine the maximum and minimum values within the window at each step. The output requires two lines: the sequence of minimums followed by the sequence of maximums.
The optimal solution employs a Monotonic Queue to achieve a time complexity of $O(N)$. Specifically, to calculate the maximums, we maintain a monotonically decreasing queue that stores array indices. As we iterate through each element in the array, we first pop the elements at the back of the queue if their corresponding values are less than or equal to the current element. This is because these smaller and older elements can never serve as the maximum for future windows. Next, the current index is pushed to the back. Then, the front of the queue is popped if its index is out of the current window scope. Finally, the value corresponding to the index at the front of the queue represents the maximum of the current window. The minimums are calculated analogously by maintaining a monotonically increasing queue. The standard solution is shown below.
Standard Solution #include<bits/stdc++.h> using namespace std; int n , a[1000005] , k ; deque<int>q ; int main() { cin >> n >> k ; for (int i = 1 ; i <= n ; i ++) { cin >> a[i] ; } for (int i = 1 ; i <= n ; i ++) { while(!q.empty() && a[i] < a[q.back()]) { q.pop_back() ; } q.push_back(i) ; if(q.front() < i - k + 1) { q.pop_front() ; } if(i >= k) cout << a[q.front()] << " " ; } cout << endl ; q.clear() ; for (int i = 1 ; i <= n ; i ++) { while(!q.empty() && a[i] > a[q.back()]) { q.pop_back() ; } q.push_back(i) ; if(q.front() < i - k + 1) { q.pop_front() ; } if(i >= k) cout << a[q.front()] << " " ; } }
Initially, this problem involves 96 Wrong Codes (WCs). After applying WrongSelect, only 8 basic WCs are retained. Their failure signatures are presented below:
$\mathcal{I}^{*}=\begin{bmatrix}0&0&0&0&0&0&0&1&1&0\\ 1&1&1&1&0&0&1&1&1&1\\ 0&0&0&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&1&0&1&1\\ 0&0&0&1&0&0&0&0&1&0\\ 0&0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&0&0&1&0&0\\ 1&1&0&0&0&0&0&0&0&0\end{bmatrix}$
C.1 Basic Wrong Codes
We meticulously analyze the retained Basic WCs and characterize their underlying error patterns.
Basic WC1 fails due to insufficient memory allocation for the queue array, where the size is set to 510,000 instead of the required 1,000,010.
[…]
Basic WC2 exhibits an incorrect order of operations where the answer is retrieved before updating the tail with the current element, causing the current element to be ignored in every window.
[…]
Distinct from the queue error in Basic WC1, Basic WC3 allocates insufficient memory for the input array using a size of $10^{5}$ rather than the required $10^{6}$.
Basic WC3 (0000001000) #include<bits/stdc++.h> using namespace std; int n,k; int tail,front; struct node { int pos,val; } q[100000010]; int a[100010]; // 100010 -> 1000010 int main() { scanf("%d%d",&n,&k); for (int i=1;i<=n;i++) { scanf("%d",&a[i]); } front=1; tail=1; q[1].val=a[1]; q[1].pos=1; for (int i=2;i<=n;i++) { while (tail>=front && q[tail].val>=a[i]) tail--; q[++tail].val=a[i]; q[tail].pos=i; while (q[tail].pos-q[front].pos+1>k) front++; if (i>=k) cout<<q[front].val<<" "; } cout<<endl; front=1; tail=1; q[1].val=a[1]; q[1].pos=1; for (int i=2;i<=n;i++) { while (tail>=front && q[tail].val<=a[i]) tail--; q[++tail].val=a[i]; q[tail].pos=i; while (q[tail].pos-q[front].pos+1>k) front++; if (i>=k) cout<<q[front].val<<" "; } }
Basic WC4 contains a subtle logic error in queue maintenance by performing an unnecessary and erroneous comparison with the head element while updating the tail. This additional operation prevents current elements from entering the queue, causing the queue to potentially become empty during the sliding process. In this state, accessing q1.top() triggers undefined behavior, retrieving residual garbage data from the underlying memory address.
[…]
Basic WC5 represents a scope error where the head and tail pointers of the queue are incorrectly re-initialized inside the loop.
[…]
Basic WC6 attempts a Sparse Table (ST) optimization but fails due to an implementation error where the allocated table size is too small for the problem constraints.
[…]
Basic WC7 attempts to fix the boundary error seen in Basic WC6 by incrementing the Sparse Table size by 1, yet it remains insufficient for the maximum constraint.
[…]
Basic WC8 incorrectly updates the head pointer instead of the tail pointer during the first element’s insertion. Additionally, it omits the insertion of the first element when initializing the second queue.
Basic WC8 (1100000000) #include<bits/stdc++.h> using namespace std; const int N=1e6+3; int n,k; int a[N]; int h=0,t=-1; int q[N]; int main() { cin>>n>>k; for (int i=1;i<=n;i++) { cin>>a[i]; } q[++h]=1; // q[++t]=1 for (int i=2;i<=n;i++) { while(i-k+1>q[h]&&h<=t)h++; while(h<=t&&a[i]<=a[q[t]])--t; q[++t]=i; if(i>=k)cout<<a[q[h]]<<" "; } cout<<endl; h=0,t=-1; // add q[++t]=1 for (int i=2;i<=n;i++) { while(i-k+1>q[h]&&h<=t)h++; while(h<=t&&a[i]>=a[q[t]])--t; q[++t]=i; if(i>=k)cout<<a[q[h]]<<" "; } return 0; }
The eight Basic WCs effectively map to the specific requirements of data structures and algorithms inherent to this problem. These error patterns include resource allocation for different variables, as well as the position, order, and conditions for queue initialization and maintenance.
On one hand, Basic WCs cover boundary constraints across different variables and granularities. For instance, Basic WC3 represents a resource error in the input array while Basic WC1, Basic WC6, and Basic WC7 target the queue. Specifically, the internal hierarchy among Basic WC1, Basic WC6, and Basic WC7 introduces a tiered validation mechanism. A less advanced test case generator, which could produce medium-scale inputs but struggles with maximum constraints, can still identify Basic WC1 and receive partial credit. This effectively avoids the “all-or-nothing” scoring trap, ensuring that the benchmark gives non-zero scores to generators that possess intermediate capabilities. Conversely, only top-tier generators that hit the absolute maximum boundary can exclude all these Basic WCs to achieve a perfect score.
On the other hand, the basis preserves logic specificities. Basic WC5 incorrectly re-initializes queue pointers inside the loop, causing the state of the sliding window to be lost at every iteration. To expose this, the test case generator must produce inputs where the window’s extremum is determined by a historical element rather than the current one, verifying the persistence of the queue. Basic WC8 exhibits dual failures, specifically on the first element’s insertion and the second queue’s initialization. This forces the generator to produce edge cases where the first element is the strict maximum or minimum for the initial windows. By retaining these basic error patterns, TC-Bench ensures that the evaluation reflects a model’s ability to cover the entire spectrum of the solution space.
C.2 Excluded Wrong Codes
Following the analysis of the Basic WCs, we proceed to examine the Excluded WCs to verify whether their error patterns are effectively encapsulated by the basis. We select Excluded WC1 as a representative example, whose failure signature is reconstructed by the combination of Basic WC8 and Basic WC4. Excluded WC1 exhibits a classic Off-by-one boundary error. During the sliding process, the code fails to timely pop the element exiting the window, causing the queue to retain invalid, expired data throughout both the initialization and maintenance phases. Crucially, this composite behavior is spanned by the basis. Basic WC8 precisely mirrors the initialization failure, as it retains stale data from the first queue due to a missing head pointer update. Meanwhile, Basic WC4 captures the maintenance failure, where additional comparison causes the queue to become empty. In this state, accessing the queue retrieves residual garbage data from memory. Together, these underlying mechanisms fully cover the error pattern of Excluded WC1.
[…]
Excluded WC1 (1101001011) #include<bits/stdc++.h> using namespace std; const int N=1000005; int a,b; int g[N],num[N],q[N],f1[N],f2[N]; int main() { scanf("%d%d",&a,&b); for (int i=1;i<=a;i++) { scanf("%d",&g[i]); } int head=1,tail=1; for (int i=1;i<=a;i++) { while(num[head]<i-b&&head<=tail) // i-b+1 head++; while(g[i]<=q[tail]&&head<=tail) tail--; num[++tail]=i; q[tail]=g[i]; f1[i]=q[head]; } head=1,tail=0; for (int i=1;i<=a;i++) { while(num[head]<i-b+1&&head<=tail) head++; while(g[i]>=q[tail]&&head<=tail) tail--; num[++tail]=i; q[tail]=g[i]; f2[i]=q[head]; } for (int i=b;i<=a;i++) cout<<f1[i]<<" "; cout<<endl; for (int i=b;i<=a;i++) cout<<f2[i]<<" "; cout<<endl; return 0; }
Similarly, we analyze Excluded WC2, whose failure signature corresponds to the combination of Basic WC8 and Basic WC5. Excluded WC2 contains a boundary error in the monotonic queue maintenance. By using the fixed condition t¿=1 instead of the dynamic h¡=t, the tail pointer can incorrectly decrement past the head pointer, violating the valid window scope and accessing invalid historical data. This error pattern is also effectively spanned by the basis. Basic WC8 captures the initialization failure, where the pointers fail to correctly establish the queue’s start (updating head instead of tail), reflecting the error’s mishandling of the absolute beginning. Basic WC5 captures the scope maintenance failure, where the queue’s dynamic state is ignored (resetting pointers inside the loop), mirroring how Excluded WC2 ignores the dynamic head boundary and corrupts the persistent state.
[…]
Excluded WC2 (1101000010) #include<iostream> #include<cstdio> #include<cstring> using namespace std; const int N=1e6+5; int n,k,a[N],q[N],p[N]; int main() { scanf("%d%d",&n,&k); for (int i=1;i<=n;i++) scanf("%d",&a[i]); int h=1,t=0; for (int i=1;i<=n;i++) { while(q[t]>a[i]&&t>=1) t--;// t>=1 -> h<=t q[++t]=a[i],p[t]=i; while(p[h]<i-k+1&&h<=t) h++; if(i>=k) printf("%d ",q[h]); } memset(q,0,sizeof(q)); memset(p,0,sizeof(p)); cout<<endl; h=1,t=0; for (int i=1;i<=n;i++) { while(q[t]<a[i]&&t>=1) t--;// t>=1 -> h<=t q[++t]=a[i],p[t]=i; while(p[h]<i-k+1&&h<=t) h++; if(i>=k) printf("%d ",q[h]); } return 0; }
C.3 Repeated Wrong Codes
Finally, we verified whether identical failure signatures indeed correspond to semantically equivalent error patterns. We selected a cluster of Repeated WCs sharing the binary signature 00010001110.
We specifically examine the representative example, Repeated WC1, shown below. This code exhibits a logic flaw during the queue maintenance phase: when inserting the current integer, the code compares it against the queue head rather than the queue tail. In a monotonic queue, the tail elements must be compared and popped to maintain monotonicity. Comparing against the head (which typically holds the window’s extremum) creates an irrelevant condition. Consequently, elements that should have been removed remain in the queue, corrupting the window’s state.
We thoroughly inspected other WCs within this same signature cluster. While they exhibit syntactic variations in implementation, we confirm that they all share the exact same root cause: the failure to correctly remove invalid elements due to flawed comparison logic. This confirms that our signature-based grouping effectively captures semantically similar faults. Due to space constraints, only key segments of these Repeated WCs are presented below.
[…]
Repeated WC 2 // Boundary offset error head = 1; tail = 0; for (int i = 1; i <= n; ++i) { while (head <= tail && i-k-1) head++;// Should be: i-k+1 while (head <= tail && a[q[tail]] <= a[i]) tail--; q[++tail] = i; if (i >= k) cout << a[q[head]] << " "; } Repeated WC 3 // Incorrectly checks queue tail (r) for expiration for (int i=k;i<=n;i++) { while(l<=r&&a[maxn[r]]<a[i]) r--; r++; maxn[r]=i; while(l<=r && maxn[r] < i-k+1) l++; // Should be: maxn[l] cout<<a[maxn[l]]<<" "; } Repeated WC 4 // Incorrectly accesses value array ’a’ instead of index array ’b’ for (int i=1;i<=n;i++) { if(hh<=tt&&a[b[hh]]<=i-k) { // Should be: b[$h$] hh++; } while(hh<=tt&&a[b[tt]]<=a[i]) { $t$--; } Repeated WC 5 // Incorrectly compares with queue front while updating tail for (int i = 1; i <= n; i++) { while (!q.empty() && a[q.front()] < a[i]) { // Should be: q.back() q.pop_back(); } q.push_back(i); if (i - q.front() >= m) { q.pop_front(); } if (i >= m) cout << a[q.front()] << " "; }
C.4 Diagnosing Realistic Scenarios
To further verify the diagnostic value of our benchmark in a realistic setting, we conducted an evaluation using the SOTA combination: Claude-4-Thinking with the LCB. We generated 40 test cases. The results showed that the generated test suite successfully excluded 6 out of the 8 Basic WCs but failed to expose Basic WC6 and Basic WC7. They are resource allocation errors requiring queue capacities of approximately $3\times 10^{5}$ and $1\times 10^{6}$, respectively. Triggering these specific faults requires forcing the monotonic queue to fill up to these limits. Mathematically, this demands a worst-case scenario where both the window size $K$ and the array length $N$ approach $10^{6}$, and crucially, the input array must follow a specific monotonic pattern (e.g., strictly increasing or decreasing) to ensure enough elements are pushed into the queue. We manually inspected all 40 generated test cases and confirmed that while the model generated large random arrays, it failed to construct this specific, structurally extreme boundary case. This demonstrates that TC-Bench effectively points out a specific weakness in current SOTA generation methods. Ultimately, this confirms that TC-Bench significantly streamlines diagnostic analysis: by narrowing the analytical scope from the entire raw dataset to a compact set of Basic WCs, it enables researchers to pinpoint model weaknesses through just a few representative examples rather than sifting through massive redundancy.
This case study provides strong empirical evidence for the practical effectiveness of our method. First, the retained Basic WCs are confirmed to be different error patterns. Second, the analysis of Excluded WCs demonstrates that redundant codes are essentially composite errors. Third, the inspection of Repeated WCs confirms that identical failure signatures reliably map to semantically equivalent root causes, validating our signature-based grouping strategy. Fourth, the real-world evaluation highlights the benchmark’s discriminative power. Collectively, these results affirm that TC-Bench successfully constructs a compact, rigorous, and representative error space, capable of delivering fine-grained and high-sensitivity evaluations for test case generation.